OpenStudy (anonymous):

did i get this correct and can you guys and help me with the question i didn't finish thx

OpenStudy (anonymous):

4. Use the following arithmetic sequence and the formula to answer the questions below. 123, 116, 109, 102, 95, ... Part I: Find the value of each of the following: a1 = _123______, a2 = __116_____, and d = ___ a2-a1 = 116-123 = -7 ____. (3 points) Part II: Find the explicit formula. Show your work. an = a1 + d(n-1) = 123 - 7(n-1) = 123 -7n + 7 = 130 - 7n a100 = 130 - 7*100 = 130 - 700 = -570 Part III: Use the explicit formula you found in Part II to find the value of the 100th term in the sequence, a100. Show your work.